OpenStudy (anonymous):

i need help with trinomials x2+5x-14

OpenStudy (smurfy14):

\[(x+7)(x-2)\]

OpenStudy (anonymous):

n2-16n-36

OpenStudy (smurfy14):

(n+2)(n-18)

OpenStudy (smurfy14):

did u ever find the button mdb?

OpenStudy (anonymous):

6k2-5k+1

OpenStudy (anonymous):

no i didnt! ive been clicking random stuff

OpenStudy (smurfy14):

(3k-1)(2k-1) ok if you see my name (smurfy14) it is in a gray box, to the end of the grey box to the right there is a blue button that says "good answer" lol

OpenStudy (anonymous):

i dont have that on there! 2y2-2y-12

OpenStudy (smurfy14):

hmm thats strange. does it not look like this?

OpenStudy (smurfy14):

2(y+2)(y-3)

OpenStudy (anonymous):

thanks :) are you good at redical expressions?

OpenStudy (smurfy14):

hmm lets see the problem

OpenStudy (anonymous):

i mean radical expressions* (in a square root) 45x2y3m7

OpenStudy (smurfy14):

\[3m^3xy \sqrt{5ym}\]

OpenStudy (anonymous):

that you so much! what about 32x6y8m12k9a22?

OpenStudy (anonymous):

thanks*

OpenStudy (anonymous):

????

OpenStudy (anonymous):

2d+3(ina sq rt)=-5

OpenStudy (smurfy14):

one sec

OpenStudy (anonymous):

ok:)

OpenStudy (smurfy14):

\[\sqrt{2d+3}=-5\] correct?

OpenStudy (anonymous):

yess!

OpenStudy (smurfy14):

2d+3=25 2d=22 d=11

OpenStudy (anonymous):

thanks :) w-2(sq rt)-3=3